OpenStudy (anonymous):

The measures of the angles of a triangle are in the extended ratio 2:8:10. What is the measure of the smallest angle? A. 18 B. 20 C. 72 D. 90

OpenStudy (anonymous):

@amistre64

OpenStudy (raden):

the smallest angle = 2/(2+8+10) * 180 = .. ?

OpenStudy (anonymous):

@RadEn 1,800?

OpenStudy (anonymous):

Is it D, by any chance?

OpenStudy (raden):

2/(2+8+10) * 180 = 2/20 * 180 = 1/10 * 180 = 180/10 = 18

OpenStudy (anonymous):

@RadEn Oh! Thank you so much for explaining that! Can you help me with some more?